OpenStudy (elise_a18):

Which statement is true about the product square root of 2(5 + square root of 8)? It is rational and equal to 7. It is rational and equal to 9. It is irrational and equal to 4 + square root of 10. It is irrational and equal to 4 + 5square root of 2.

OpenStudy (elise_a18):

@sammixboo

OpenStudy (trojanpoem):

I hardly believe that It's readable rootof 2(5+ rootof(8)) as complicated as magic.

OpenStudy (zarkon):

this??? \[\sqrt{2}\cdot (5+\sqrt{8})\]

OpenStudy (elise_a18):

mhm @zarkon